OpenStudy (shae.mccoy):

The paper water cups provided near the water jug are cones of radius 1.5cm and height 2cm. Your water bottle is a cylinder of radius 1.5cm and height 4 cm. How many full paper cups of water would you have to drink to get the same amount of water as one full water bottle? @cwrw238 @iGreen @ikram002p @perl

OpenStudy (shae.mccoy):

I give medals

OpenStudy (shae.mccoy):

@perl

OpenStudy (shae.mccoy):

@DanielleJayla

OpenStudy (igreen):

Find the volumes of the Paper Cups and the Water bottle. Volume of a Cone: \(\dfrac{1}{3}\pi r^2h\) Volume of a Cylinder: \(\pi r^2 h\)

OpenStudy (igreen):

Let's find the volume of the cone first: \(V = \dfrac{1}{3}\pi r^2 h\) \(V =\dfrac{1}{3}(3.14)(1.5^2)(2)\) Simplify exponent: \(V =\dfrac{1}{3}(3.14)(2.25)(2)\) Multiply: \(V =(1.05)(2.25)(2)\) Multiply: \(V =(2.36)(2)\) Multiply: \(V = 4.72\) So the volume of the cone cup is 4.72.

OpenStudy (igreen):

Now let's find the Volume of the Cylinder: \(V = \pi r^2 h\) \(V = (3.14)(1.5^2)(4)\) Simplify Exponent: \(V = (3.14)(2.25)(4)\) Multiply: \(V = (7.07)(4)\) Multiply: \(V = 28.28\)

OpenStudy (igreen):

Now to find how many cups of water we need, we can just divide: 28.28 / 4.72 = ? @Sha

OpenStudy (igreen):

Can you divide 28.28 and 4.72? @Shae.McCoy

OpenStudy (shae.mccoy):

sorry.. Kinda zoned out

OpenStudy (shae.mccoy):

5.99? @ig